Jane and John Smith have accumulated $650,000 for their retirement, which begins today. They plan to receive monthly payments from their investments, which will be paid at the beginning of each month over the next 35 years based on their estimated life expectancy. If investments are accumulating at an after-tax annual rate of 5.75%, compounded monthly, what will be the payment amount that the Smiths will receive each month to the nearest dollar?

Please be very thorough with work and provide specific formula used.
Answer #1

PV of annuity due = Annuity + Annuity*(1-1/(1+rate)^(number of terms-1))/rate

Here

PV = 650000

Rate = 5.75%/12

Number of terms = 35*12


650000 = A+ A*(1-1/(1+0.0575/12)^(35*12-1))/ (0.0575/12)

650000 = A+ 180.53394A

A = 650000/181.53394

=$3580.60

Amount received per month = $3580.60